A spy report about a suspected car reads as follows. “The car moved 2∙00 km towards east, made a perpendicular left turn, ran for 500m, made a perpendicular right turn, ran for 4∙00 km and stopped”. Find the displacement of the car.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. (AD) ⃗ 2i ̂ + 0.5 j ̂ + 4 K ̂ = 6i ̂ + 0.5 j ̂ AD = √(AE^2+DE^2 ) = 6.02 KM Tan θ = DE / AE = 1/12 Θ = tan 1 (1/12) The displacement of the car is 6.02 km along the distance tan-1 (1/12) with positive x-axis.